OpenStudy (anonymous):

The number of cars at a dealership dropped from 64 to 48 after a weekend sale. What is the percent of decrease in the number of cars? A. 20% B. 25% C. 33% D. 75%

OpenStudy (catlover5925):

take 48 and divide it by the bigger # its like if you want to know ur % grade on a test you take how many you got right and divide it by the total #

OpenStudy (anonymous):

\[64 - 48 = 16\]

OpenStudy (jjuden):

b

OpenStudy (catlover5925):

so you would get 0.75 and move the decimal over and it would be 75%

OpenStudy (anonymous):

\[16 \div 24 = .25 = 25%\]

OpenStudy (anonymous):

so its not b? @Catlover5925

OpenStudy (catlover5925):

@mg_omg it is not 25%

OpenStudy (jjuden):

its 25%

OpenStudy (catlover5925):

no it is D. 75%

OpenStudy (anonymous):

What is the percent of decrease from 320 to 32? A. 10% B. 90% C. 100% D. 288% @Catlover5925 i got b, is that correct?

OpenStudy (anonymous):

im confused!

OpenStudy (catlover5925):

meee toooooooooooooo idk what I did wrong, tag some mods in they can help

OpenStudy (jjuden):

yes let me prove it hold on

OpenStudy (anonymous):

it's 25%

OpenStudy (anonymous):

the percentage of decrease from 64 to 48 is 25%

OpenStudy (jjuden):

\[\frac{ 64-48 }{ 64 }* 100=\frac{ 16 }{ 64 }* 100\] \[=\frac{ 1 }{ 4 }*100=25%\]

OpenStudy (anonymous):

^

OpenStudy (jjuden):

25% is thy answer

OpenStudy (anonymous):

well...that seems right so ill enter that answer! thank you all for helping me

OpenStudy (jjuden):

welcome